# Building a Balanced k-d Tree in O(kn log n) Time

### v43

@@ 5 Conclusion
The source code that was provided with the original version of this article [7] and that implemented the $O\left(n\log n\right)$ and $O\left(kn\log n\right)$ k-d tree-building algorithms is superseded by the optimized source code described in Section 6.9.
@@ 6.2 Optimized Merge Sort
Both tree-building algorithms depend on merge sort [16] that copies elements between arrays. An optimized merge sort minimizes copying via four variants of merge sort [17].
@@ 6.10 Least-Squares Curve Fitting
However, for a number of threads $q>2$, the average values $\hat{f}_{i}^{q}$ and $\hat{g}_{i}^{q}$ are calculated for $f_{i}$ and $g_{i}$ respectively because the number of threads $q$ is not constant across all levels of the nascent k-d tree, as illustrated by the following example.

### v44

@@ 5 Conclusion
Source code that implemented the $O\left(n\log n\right)$ and $O\left(kn\log n\right)$ k-d tree-building algorithms and that was provided with prior versions of this article has been superseded by the source code described in Section 6.9.
@@ 6.2 Optimized Merge Sort
Both tree-building algorithms depend on merge sort that copies elements between arrays [8]. An optimized merge sort minimizes copying via four variants of merge sort [15].
@@ 6.10 Least-Squares Curve Fitting
The correlation coefficient $r_{\mathrm{fgt}}$ for the least-squares fit is a coefficient of linear multiple correlation that is calculated from the linear correlation coefficients $r_{\mathrm{ft}}$, $r_{\mathrm{gt}}$, and $r_{\mathrm{fg}}$ as follows [16].
$\ r_{\mathrm{ft}}=\frac{n\sum{f_{i}t_{i}}-\sum f_{i}\sum t_{i}}{\sqrt{\left[n\sum{f_{i}}^{2}-\left(\sum f_{i}\right)^{2}\right]\left[n\sum{t_{i}}^{2}-\left(\sum t_{i}\right)^{2}\right]}}\\ \\ r_{\mathrm{gt}}=\frac{n\sum{g_{i}t_{i}}-\sum g_{i}\sum t_{i}}{\sqrt{\left[n\sum{g_{i}}^{2}-\left(\sum g_{i}\right)^{2}\right]\left[n\sum{t_{i}}^{2}-\left(\sum t_{i}\right)^{2}\right]}}\\ \\ r_{\mathrm{fg}}=\frac{n\sum{f_{i}g_{i}}-\sum f_{i}\sum g_{i}}{\sqrt{\left[n\sum{f_{i}}^{2}-\left(\sum f_{i}\right)^{2}\right]\left[n\sum{g_{i}}^{2}-\left(\sum g_{i}\right)^{2}\right]}}\\ \\ r_{\mathrm{fgt}}=\frac{r_{\mathrm{ft}}^{2}+r_{\mathrm{gt}}^{2}-2r_{\mathrm{ft}}r_{\mathrm{gt}}r_{\mathrm{fg}}}{1-r_{\mathrm{fg}}^{2}}\end{matrix}$ (10)
For a number of threads $q>2$, Equations 9 and 10 use the average values $\hat{f}_{i}^{q}$ and $\hat{g}_{i}^{q}$ respectively in place of $f_{i}$ and $g_{i}$ because the number of threads $q$ is not constant across all levels of the nascent k-d tree, as illustrated by the following example.